Compute the value of 9x²+ 4y² if xy = 6 and 3x + 2y = 12. *
1

Grade:9

2 Answers

Harshit Singh
3 years ago
Dear Student
Given That
3x + 2y = 12
On squaring both sides
9x2 + 4y2 + 12xy=144
putting xy=6
9x2 + 4 y2 = 144 –72
= 72

Thanks
Ram Kushwah
3 years ago
Given that xy=6 
and 3x+2y=12
 
Squaring both sides
(3x+2y)²=12²
(3x)²+(2y)²+2(3x)(2y)=144
9x²+4y²+12xy=144
putting xy=6
 
9x²+4y² +12*6=144
9x²+4y² =144-72
 
9x²+4y² =72